Calculate the resonant frequency of a circuit containing an inductor of 0.2 mH and a capacitor of 2F capacitance?

Grade:12th pass

1 Answers

Aman
5 years ago
Formula of resonant frequency is f=(1/2×3.14×(LC)^1/2
f=1/2×3.14×(LC)^1/2
=1/6.28×(2×2/10000)^1/2
1/6.28×2/100
=100/12.46
8.025